Montreal Seating Co., a manufacturer of chairs, had the following data for 2020:

Sales 3,200 units
Sales price $60 per unit
Variable costs $30 per unit
Fixed costs $49,500

-Calculate the contribution margin ratio.(%)

-Calculate the break-even point in dollars.

-Calculate the margin of safety in dollars.

-The company wishes to increase its total dollar contribution margin by 56% in 2021. Determine by how much it will need to increase its sales if all other factors remain constant.
Answer #1

Ans:

1.Contribution margin=(SP-VC)

=(60-30)=$30/unit

hence Contribution margin ratio=Contribution margin/SP

=(30/60)

=50%.

2.Breakeven point=Fixed costs/Contribution margin ratio

=(49500/0.5)

=$99000

3.Margin of safety=Total sales-Breakeven sales

=(3200*60)-99000

=$93,000

4.Margin of safety ratio=($51250/Total sales)

=(93000/(3200*60))

=48.44%.

5.Total contribution margin required=($30*3200)*1.56

=$149,760

Contribution margin/unit=$30

Hence number of units to be sold=($149,760/$30)=4992.

Hence Total increase in units required=(4992-3200)=1792 units

ie =(1792*60)=$107,520


Hope this helped ! Let me know in case of any queries.